a) how many hours needed to complete filling process of 150 000 mL in 1 mL bottles

*speed: 2 bottles per second


b) how many staff are needed at one time to complete packing the bottles in carton within 24 hours

* 1 carton = 10 bottles

* speed: 1 carton per 15 seconds


1
Expert's answer
2021-07-09T11:23:58-0400

a) The machine caps 2 bottles per second.


"1\ s=(1/3600)\ hr"

"(1/3600)\ hr \ \ -\ \ 2\ bottles"

"x\ hr \ \ -\ \ 150000\ bottles"

"x=\dfrac{\dfrac{150000}{3600}}{2}\ hr=20\dfrac{5}{6}\ hr"

b)


"(15/3600)\ hr \ \ -\ \ 10\ bottles"

"x\ hr \ \ -\ \ 150000\ bottles"

"x=\dfrac{\dfrac{15}{3600}(150000)}{10}\ hr=62.5\ hr"

"2<62.5\/24<3"

3 staff needed at one time.